# Efficient Online Inference of Vision Transformers by Training-Free Tokenization

### v1

@@ 4 Experiments
Figure 3: Visualization of the vocabulary of ${\mathcal{T}}_{inter}^{100}$. Patches from ImageNet-1K are matched to the closest visual word using the Euclidean distance. Visual words are shown to depict basic features such as colors or edges. As centroids, each visual word is an average representation of patches that belong to the matched cluster.
Baselines.
For image classification, we load the pre-trained CLIP [43] model from HuggingFace [^4] https://huggingface.co/openai/clip-vit-base-patch16. The prefix ”a photo of a ” is also added to each text category. An image size of 224 $\times$ 224 is used with bilinear interpolation for CLIP. For image captioning, we load the pre-trained BLIP [28] model from HuggingFace [^5] https://huggingface.co/Salesforce/blip-image-captioning-base. To perform zero-shot captioning, we use a beam size of 3 along with maximum and minimum lengths of 20 and 5, respectively. Unlike [28], we do not include the prefix ”a picture of ” as we found it to work better. An image size of 384 $\times$ 384 is used with bicubic interpolation. Both CLIP and BLIP utilize the ViT-B/16 image encoder unless stated otherwise.
@@
Aside from the pre-trained model which we denote as $Base$, we also consider 8-bit quantization [11] and token merging [5] as additional baselines. We denote the former as $Q_{8}$ and the latter as $ToME$. Following [5], we utilize a reduction per layer for $ToME$ of 13 with CLIP and 23 with BLIP due to their respective input image sizes. For the VWTs, we set the top-k of the intra-image approach to 50% of the total number of patches which we denote as ${\mathcal{T}}_{intra}^{0.5}$. For the inter-image approach, we set the threshold to 0.1 unless stated otherwise and denote it as ${\mathcal{T}}_{inter}^{{\mathcal{V}}}$, where ${\mathcal{V}}$ is the size of the vocabulary. Lastly, our experiments are conducted using a single NVIDIA A100 GPU.
@@
Figure 4: Visualization of long-form captioning on COCO. Longer captions are generated via a length penalty of 2.0 and maximum length of 40. Interestingly, the smaller vocabulary of ${\mathcal{T}}_{inter}^{100}$ improves descriptiveness and coherence more than ${\mathcal{T}}_{intra}^{0.5}$ or ${\mathcal{T}}_{inter}^{1000}$.
Inference Efficiency.
@@
First we analyze the effects of VWTs on token length. We seek to understand how the choice of pre-processing data and vocabulary size affects the degree of compression. Table 1 shows the token length per sample (including [CLS]) on different datasets. Unlike ${\mathcal{T}}_{intra}^{0.5}$, the sequence lengths induced by ${\mathcal{T}}_{inter}^{{\mathcal{V}}}$ are not equal. First, we compare ${\mathcal{T}}_{inter}^{{\mathcal{V}}}$ pre-processed on the in-domain dataset and ImageNet-1K [10]. The in-domain dataset is represented by the training split if available. On text data, in-domain tokenizers [9, 61, 40] have been shown to produce shorter sequences by specializing the vocabulary on the given distribution. Interestingly, we observe no such effect with image data as seen by the similar lengths between In-Domain and ImageNet-1K. Only on CelebA, do we see a greater reduction with ${\mathcal{T}}_{inter}^{1000}$ and ${\mathcal{T}}_{inter}^{10000}$ pre-processed on ImageNet-1K. Second, unlike text tokenizers, a trend of decreasing compression is seen as vocabulary size increases. With text, larger vocabularies ensure that more tokens are kept as words rather than subwords. In our case, we posit that an increasing number of patches are matched to separate visual words, thus lowering the overall compression. We restrict our remaining analyses to ${\mathcal{T}}_{inter}^{1000}$ pre-processed on ImageNet-1K unless stated otherwise.
@@
Having analyzed the effects on token length, we turn our attention to the practical metrics of wattage and runtime. For online inference, savings in wattage should not incur a significant cost to runtime. In Table 2, we compare the efficiency of tokenization to $Base$, $Q_{8}$, and $ToME$ on the datasets. First, we find the wattage of VWTs to be lower than $Base$ across the datasets. On CelebA, this reduction is up to 19% with ${\mathcal{T}}_{inter}^{1000}$. Naturally, efficiency decreases as vocabulary size increases due to smaller compression. Although $Q_{8}$ and $ToME$ result in a lower wattage than VWTs, we note the drawbacks in runtime. In particular, $Q_{8}$ results in a significantly longer runtime for $Base$ as tensors need to be repeatedly quantized and dequantized. Meanwhile, $ToME$ can increase runtime by up to 2×2$\times$2 × on Waterbirds, CelebA, and MetaShift. We observe increases of up to 20% at most with ${\mathcal{T}}_{inter}^{1000}$ on the aforementioned datasets. ${\mathcal{T}}_{intra}^{0.5}$ displays lower runtimes than $Base$ across the datasets while ${\mathcal{T}}_{inter}^{1000}$ does so on COCO and NoCaps. We have shown how VWTs are more suitable for online inference by striking a good balance between savings in wattage and costs to runtime.
Performance Retention of Classification, Robustness, and Captioning.
Another important factor in compression is the effect on model performance. In Table 3, we tabulate the performance of image classification and captioning using CLIP [43] and BLIP [28], respectively. For classification and group robustness, we analyze the average and worst-group (robust) performance as done by [47, 46]. First, we find the degradation in average accuracy to be small for ${\mathcal{T}}_{intra}^{0.5}$ and ${\mathcal{T}}_{inter}^{1000}$. The largest drop of up to 2% only is shown by the former on MetaShift. Likewise, $Q_{8}$ and $ToME$ retain a high average accuracy across the datasets. Second, we observe possible improvements in robustness with VWTs. On Waterbirds and CelebA, the worst-group accuracy with ${\mathcal{T}}_{intra}^{0.5}$ increases by up to 29% and 8%, respectively. Only on MetaShift is robustness lower than $Q_{8}$ or $ToME$ relative to $Base$. Like [20], we find compression to not always be harmful to group robustness.
We conduct more evaluations on the large-scale OpenImages v6 dataset for zero-shot classification. We report mean Average Precision (mAP) as done by [23]. From Tables 2 and 3, we find VWTs to be more suitable for online inference than $Q_{8}$ or $ToME$ in the large-scale setting, while retaining comparable performance to the $Base$ model.
For image captioning with BLIP, we evaluate our models following the setting in  [28] by using the BLEU, CIDEr, and SPICE metrics w.r.t. the ground truth captions. On COCO, VWTs display a higher performance than $ToME$ and are competitive with $Q_{8}$. On NoCaps, we see the largest degradation on the out-of-domain samples with ${\mathcal{T}}_{inter}^{1000}$. However, overall performance is still higher than $ToME$. Only $Q_{8}$ displays a slight improvement over $Base$. Like $ToME$, VWTs are shown to not require additional end-to-end fine-tuning for performance retention.
(a) Wattage (watt) $\downarrow$
(b) Runtime (millisecond) $\downarrow$
[…]
Visual Words.
To better understand the inter-image approach, we visualize the patch matching of ${\mathcal{T}}_{inter}^{100}$ in Figure 1. We highlight in identical colors patches that are matched with one another. First, we find that patches representing the background are more frequently grouped than those of the foreground. On Waterbirds, the merging of background patches may explain the improved robustness in Table 3 by mitigating the spurious correlation with the background. On CelebA, ${\mathcal{T}}_{inter}^{100}$ tends to avoid matching the eyes or mouths of the individuals. Second, we observe that patch matching is capable of grouping similar but non-adjacent visual concepts. In certain examples of MetaShift and NoCaps, multiple cats and foods are seemingly matched together, respectively. Our analysis shows that patch matching serves as a rudimentary form of image segmentation.
In Figure 3, we visualize the vocabulary of ${\mathcal{T}}_{inter}^{100}$ to analyze the formation of the visual words. We show patches from ImageNet-1K (i.e. pre-processing data) that are matched to each visual word using the Euclidean distance. Since visual words are centroids, patches that are matched to the same visual word belong to the same cluster. We observe that visual words depict basic features such as colors or edges. These features are also formed as an average representation of the patches that belong to each cluster. By representing basic features, visual words serve as effective anchor points to which patches can be matched to. Analysis on the sparsity of the vocabularies in the inter-image approach can be found in Section B.5.
Long-form Captioning.
The generated captions by BLIP [28] can be used as priors for further training of other models. As such, longer descriptive captions may be more desirable than the typical short captions associated with Internet data. In Figure 4, we visualize the long-form captions on COCO. To enable longer generations, we set the length penalty to 2.0 and double the maximum length to 40. All other generation settings are kept the same. With longer captions, the generation may degenerate into unnecessary repetitions on certain samples. Interestingly, descriptiveness and coherence improves more with ${\mathcal{T}}_{inter}^{100}$ than ${\mathcal{T}}_{intra}^{0.5}$ or ${\mathcal{T}}_{inter}^{1000}$. Additional long-form captions by $Q_{8}$, $ToME$, and ${\mathcal{T}}_{inter}^{10000}$ are provided in Section C.1.
Batch Processing.
Earlier we stated how VWTs works optimally for online inference. In Figure 5, we seek to better understand its effectiveness for offline inference where batch sizes are greater than 1. Using typical batches of $\{4,8,16,32\}$, we compare ${\mathcal{T}}_{intra}^{0.5}$ and ${\mathcal{T}}_{inter}^{1000}$ to $Base$ and $ToME$. First, we find that as batch size increases, lower wattages relative to $Base$ are seen with ${\mathcal{T}}_{intra}^{0.5}$ and $ToME$, particularly at lower batch sizes. Both VWTs and $ToME$ eventually reach wattage parity with $Base$ at larger batches (i.e. 32), with ${\mathcal{T}}_{inter}^{1000}$ doing so much earlier. Second, we observe different runtime trends between VWTs and $ToME$. As batches grow, ${\mathcal{T}}_{inter}^{1000}$ continues to show similar runtimes to $Base$. Meanwhile, ${\mathcal{T}}_{intra}^{0.5}$ and $ToME$ display lower runtimes with the exception of batch size 4 for the latter. As such, during batching, the inter-image approach does not confer the same degree of efficiency improvements.
When to use the intra-image or inter-image approach?
Concerning the choice between the intra-image or inter-image approach, we posit that when global information is required, dropping tokens may be more advantageous by removing unnecessary noise from the visual input. On the other hand, when the task necessitates local information (e.g. long-form captioning, object detection, etc.), merging tokens may better preserve the visual concepts. For example, the image on the first row of Figure 7 shows that ${\mathcal{T}}_{intra}^{0.5}$ removes the mountainous background, thus leading to the absence of ”mountain” in the long-form caption of Figure 4. Potentially, combining both approaches may maximize compression while preserving information by merging the mountains and dropping the bushes in said image.
5 Conclusion
In this work, we set out to define a training-free tokenization for ViTs that lowers wattage while balancing costs to runtime and performance. In online scenarios, we have shown empirically that our intra-image and inter-image approaches are competitive with 8-bit quantization and token merging for image classification and captioning. Qualitatively, we observe how the inter-image approach groups analogous visual concepts via visual words that represent basic features. Analysis on large-scale classification further validates the viability of our method while long-form captioning shows its potential for improving descriptiveness and coherence. Lastly, the inter-image approach has limits for batch processing due to its non-parallelism. As a future work, further research can be made into combining both approaches.
Acknowledgements
Appendix A Further Details
A.1 Datasets for Group Robustness
Here, we detail the task of each dataset for group robustness. In Table 4, we also tabulate the labels and attributes that define each subgroup along with their sample sizes.
Waterbirds.
Given an image of a bird, the task is to predict whether it is a waterbird or landbird [57]. The attribute is the background that the bird is on [47].
CelebA.
Given an image of a person, the task is to predict whether their hair color is blond or not [33]. The attribute is the binary gender of the person.
MetaShift.
Given an image of an animal, the task is to predict whether it is a dog or cat [29]. The attribute is the environment that the dog or cat is in.
Dataset Subgroup Label Attribute Training Validation Test
Waterbirds 0 0 (landbird) 0 (on land) 3498 467 2255
1 0 (landbird) 1 (on water) 184 466 2255
2 1 (waterbird) 0 (on land) 56 133 642
3 1 (waterbird) 1 (on water) 1057 133 642
CelebA 0 0 (non-blond) 0 (woman) 71629 8535 9767
1 0 (non-blond) 1 (man) 66874 8276 7535
2 1 (blond) 0 (woman) 22880 2874 2480
3 1 (blond) 1 (man) 1387 182 180
MetaShift 0 0 (dog) 0 (outdoor) 784 127 273
1 0 (dog) 1 (indoor) 507 75 191
2 1 (cat) 0 (outdoor) 196 33 65
3 1 (cat) 1 (indoor) 789 114 345
Table 4: Defined subgroups in Waterbirds, CelebA, and MetaShift.
Appendix B Supplementary Experiments
Here, we detail the supplementary experiments that we conducted. First, we explore additional design choices and effects of the inter-image approach. Section B.1 shows the ineffectiveness of patch matching with visual words formed in the embedding space. In Section B.2, random matching is applied to further determine if matching by the visual words are meaningful. Section B.3 ablates the similarity threshold by relaxing it for further compression. The fairness of tokenization and sparsity of the vocabulary are shown in Sections B.4 and B.5, respectively.
B.1 Ineffectiveness of the Embedding Space
In Table 5, we analyze the sequence compression using visual words formed in the embedding space. To do so, we initialize each vocabulary by clustering embeddings from the pre-trained convolution layer of CLIP or BLIP. During inference, we match the activations from the pre-trained convolution layer with the visual words. Compared to Table 1, we observe a notable reduction in the degree of compression irrespective of pre-processing data and vocabulary size. In [50], embeddings are shown to become progressively more similar to one another due to self-attention. Hence, it is unsurprising that matching activations and visual words at the beginning of the network is ineffective.
B.2 Random Matching of Visual Words
In Table 6, we study the effects of randomly matching the patches. We initialize the pairwise cosine distance by sampling from a uniform distribution of $[0,2]$. First, we find the token length (including [CLS]) to differ noticeably from Table 1. For ${\mathcal{T}}_{inter}^{100}$, sequences are further reduced across the datasets. Conversely, ${\mathcal{T}}_{inter}^{1000}$ and ${\mathcal{T}}_{inter}^{10000}$ display little compression. Second, performance is shown to change from Table 3. We observe a significant degradation in average and worst-group accuracies with ${\mathcal{T}}_{inter}^{100}$ on Waterbirds and MetaShift. For ${\mathcal{T}}_{inter}^{1000}$ and ${\mathcal{T}}_{inter}^{10000}$, performance does not shift much from $Base$ as barely any compression occurs. With random matching, the captions become completely different than those of $Base$ in Figure 10.
B.3 Ablation of the Similarity Threshold
We have shown how VWTs can improve the efficiency of online inference. To better understand their limitations, we ablate the threshold of ${\mathcal{T}}_{inter}^{100}$ by setting it to values of $\{0.2,0.3,0.4,0.5\}$ in Table 7. We seek to determine if exploiting higher thresholds for increased compression is a viable strategy. Naturally, we observe a reduction in performance as increasingly dissimilar patches are merged. For Waterbirds and MetaShift, the worst-group accuracy degrades more significantly than the average, especially with the former. Interestingly, average accuracy remains relatively unchanged while worst-group accuracy improves significantly on CelebA irrespective of similarity threshold. We posit that at higher thresholds, the merging of core features represented by the foreground object results in the reduced performance of Waterbirds and MetaShift.
B.4 Fairness of the Tokenization
On text data, tokenizers have been shown to induce unfairness between languages [41, 2]. Since most LLMs are built using English data, tokenization of minority languages produces inherently longer sequences that raises compute costs. We seek to analyze if similar effects exist with VWTs as well. In Table 8, we show the breakdown in token length (including [CLS]) and accuracy (w.r.t $Base$) by subgroup. First, we observe a notable difference in compression between the subgroups of Waterbirds. With ${\mathcal{T}}_{100}$, sequences might differ by up to 39 tokens as seen with subgroups 0 and 3. Smaller discrepancies are displayed on CelebA and MetaShift except for ${\mathcal{T}}_{100}$ on the former. Second, we find that compression does not affect all subgroups equally. Accuracy improves on certain subgroups and degrades on others. A stronger compression does not correlate also with a large change in performance. Like text tokenizers, VWTs may induce inequality in compression that leads to varying inference costs and performance for different subgroups.
B.5 Sparsity of the Vocabulary
To better understand the utilization of the visual words, we plot the probability distribution of the matches in Figure 6. Regardless of the dataset, we find that certain visual words are matched more frequently than others, thus leading to a large skew in the distributions. Greater sparsity is also displayed by larger vocabularies as many visual words remain unused across datasets. As such, pruning may be applied to achieve a more efficient vocabulary size post-clustering.
Appendix C Additional Results
Here, we provide additional results by other models and hyperparameters in our experiments. Section C.1 lists long-form captions by $Q_{8}$, $ToME$, and ${\mathcal{T}}_{inter}^{10000}$. Sections C.2 and C.3 detail the intra-image and inter-image approaches with various dropping ratios and vocabulary sizes, respectively.
C.1 Long-form Captions
In Figure 7, we showcase long-form captions by $Q_{8}$, $ToME$, and ${\mathcal{T}}_{inter}^{10000}$. First, we observe improvements in the accuracy of the descriptions with $Q_{8}$ over $Base$ in Figure 4. Although the captions relate better to the image, the overall length is shorter than those of other methods. The generation may continue to introduce unnecessary repetitions. Second, we find $ToME$ and ${\mathcal{T}}_{inter}^{10000}$ to also be beneficial for descriptiveness and coherence. On certain samples, repeating n-grams may still occur.
C.2 Dropping Ratio
Following [39], we apply the intra-image approach with ratios of $\{0.25,0.33,0.7\}$ to the total number of patches in Table 9. We observe a natural degradation as the ratio increases with performance dropping steeply at 0.7. In Figures 8 and 9, we visualize the dropped patches and captions on COCO, respectively. In most cases, the patches with the lowest pixel variance correspond to uninformative backgrounds. Dropping such patches may continue to preserve the foreground object.
C.3 Vocabulary Size
In Table 10, we apply the inter-image approach with vocabulary sizes of $\{100,10000\}$. We find performance to be noticeably lower than $Base$ in Table 3 with the smaller vocabulary of ${\mathcal{T}}_{inter}^{100}$. Meanwhile, performance retention is higher for ${\mathcal{T}}_{inter}^{10000}$ due to lower sequence compression in Table 1. When visualizing the captions on COCO in Figure 10, we also observe that increasing thresholds cause larger deviations from $Base$ than reducing the vocabulary size.
Dataset ${\mathcal{T}}_{inter}^{100}$ ${\mathcal{T}}_{inter}^{1000}$ ${\mathcal{T}}_{inter}^{10000}$
@@ C.3 Vocabulary Size
COCO 405 406 409 408 442 440
@@
Table 5: Token length per sample (including [CLS]). Tokenizers of varying pre-processing data and vocabulary sizes are shown. Visual words are formed from embeddings in the pre-trained convolution layer of CLIP or BLIP and not the image patches. Poor compression is seen irrespective of pre-processing data and vocabulary size due to the lack of similarity between the activations and visual words.
Dataset Model ${\mathcal{T}}_{intra}^{100}$ ${\mathcal{T}}_{intra}^{1000}$ ${\mathcal{T}}_{intra}^{10000}$

### v2

@@ 4 Experiments
Figure 4: Visualization of the vocabulary of ${\mathcal{T}}_{inter}^{100}$. Patches from ImageNet-1K are matched to the closest visual word using the Euclidean distance. Visual words are shown to depict basic features such as colors or edges. Each visual word is an average representation of patches that belong to the matched cluster.
4.1 Datasets and Settings
We conduct our analysis through the lens of (i) classification performance of visual recognition, (ii) subgroup robustness, and (iii) generative performance of visual captioning. For (i) and (ii), we utilize three publicly available datasets (Waterbirds [63], CelebA [36], MetaShift [31]) that are typical benchmarks in robustness and fairness research [50, 35, 70]. To perform (zero-shot) classification, we compute the cosine similarity between an image embedding and the following encoded text labels4 for Waterbirds, CelebA, and MetaShift, respectively: {’landbird’, ’waterbird’}, {’non-blond’, ’blond’}, {’dog’, ’cat’}. Further details on the defined subgroups are provided in Appendix 6.1. For (i), we also conduct a large-scale evaluation on the OpenImages v6 dataset [29]. Following Huang et al. [25], the test split is divided into common and rare subsets that consist of 57 224 and 21 991 images, respectively. The former has 214 classes, while the latter has 200. To perform (zero-shot) classification, we compute the cosine similarity between an image embedding and the encoded text label [^4] The prefix ”a photo of a ” is also added to encode each text label.. For (iii), we utilize the Karpathy test split of COCO dataset [33] and a validation set of NoCaps dataset [1] following the setting in previous work [30]. Finally, to study inference efficiency, we utilize all datasets for the visual tasks (i)-(iii) described above. We compute the efficiency per sample and average across all samples. For the wattage calculation, we call ”pynvml.nvmlDeviceGetPowerUsage(handle)/1000” as done by Weights and Biases, which leverages the NVIDIA Management Library for monitoring and managing the NVIDIA GPUs. The experiments are also conducted in an isolated environment, thus ensuring the efficiency measurements are as accurate as possible. Note that we do not include the Pre-Process step of the inter-image approach into our efficiency calculations as this process is only done once and may be reused multiple times.
4.2 Implementation Details
@@
For image classification, we load the pre-trained CLIP [46] model from HuggingFace [^5] https://huggingface.co/openai/clip-vit-base-patch16. An image size of 224 $\times$ 224 is used with bilinear interpolation for CLIP. For image captioning, we load the pre-trained BLIP [30] model from HuggingFace [^6] https://huggingface.co/Salesforce/blip-image-captioning-base. To perform zero-shot captioning, we use a beam size of 3 along with maximum and minimum lengths of 20 and 5, respectively. An image size of 384 $\times$ 384 is used with bicubic interpolation. Both CLIP and BLIP utilize the ViT-B/16 image encoder unless stated otherwise. Aside from the pre-trained model which we denote as $Base$, we also consider 8-bit quantization [11] and token merging [5] as additional baselines. We denote the former as $Q_{8}$ and the latter as $ToME$. Following Bolya et al. [5], we utilize a reduction per layer for $ToME$ of 13 with CLIP and 23 with BLIP due to their respective input image sizes. For the VWTs, we set the top-k of the intra-image approach to 50% of the total number of patches which we denote as ${\mathcal{T}}_{intra}^{0.5}$. For the inter-image approach, we set the threshold to 0.1 unless stated otherwise and denote it as ${\mathcal{T}}_{inter}^{{\mathcal{V}}}$, where ${\mathcal{V}}$ is the size of the vocabulary. Lastly, our experiments are conducted using a single NVIDIA A100 GPU. Since our focus is on the online setting (real-time), we set the batch size to 1 unless stated otherwise.
@@
Figure 5: Visualization of long-form captioning on COCO. Longer captions are generated via a length penalty of 2.0 and a maximum length of 40. Interestingly, the smaller vocabulary of ${\mathcal{T}}_{inter}^{100}$ possesses higher descriptiveness and coherence than ${\mathcal{T}}_{intra}^{0.5}$, ${\mathcal{T}}_{inter}^{1000}$, or ${\mathcal{T}}_{inter}^{10000}$ despite its lesser compression.
4.3 VWTs and Inference Efficiency
@@
First we analyze the effects of VWTs on token length. We seek to understand how the choice of pre-processing data and vocabulary size affects the degree of compression. Table 1 shows the token length per sample (including [CLS]) on different datasets. Unlike ${\mathcal{T}}_{intra}^{0.5}$, the sequence lengths induced by ${\mathcal{T}}_{inter}^{{\mathcal{V}}}$ are not equal. First, we compare ${\mathcal{T}}_{inter}^{{\mathcal{V}}}$ pre-processed on the in-domain dataset and ImageNet-1K [10]. The in-domain dataset is represented by the training split if available. On text data, in-domain tokenizers [18, 19, 9, 67, 43] have been shown to produce shorter sequences by specializing the vocabulary on the given distribution. Interestingly, we observe no such effect with image data as seen by the similar lengths between In-Domain and ImageNet-1K. Only on CelebA, do we see a slightly greater reduction with ${\mathcal{T}}_{inter}^{1000}$ and ${\mathcal{T}}_{inter}^{10000}$ pre-processed on ImageNet-1K. Second, unlike text tokenizers, decreasing compression is seen as vocabulary size increases. With text, larger vocabularies ensure that more tokens are kept as words rather than subwords. We posit that an increasing number of patches are matched to separate visual words, thus lowering the overall compression.
@@
Having analyzed the effects on token length, we turn to the practical metrics of wattage and runtime. For online inference, savings in wattage should not incur a significant cost to runtime. In Table 2, we compare the efficiency of VWTs to $Base$, $Q_{8}$, and $ToME$. Note that we compute both metrics using the image encoder of CLIP or BLIP only. First, we find the wattage of VWTs to be lower than $Base$ across the datasets. On COCO, this reduction is up to 25% with ${\mathcal{T}}_{inter}^{100}$. Naturally, efficiency decreases as vocabulary size increases due to smaller compression. Although $Q_{8}$ and $ToME$ result in a lower wattage than VWTs, we note the drawbacks in runtime. In particular, $Q_{8}$ results in a significantly longer runtime for $Base$ as tensors need to be repeatedly quantized and dequantized. Meanwhile, $ToME$ can increase runtime by up to 100% on Waterbirds, CelebA, MetaShift, and OpenImages. We observe increases of up to 20% at most with ${\mathcal{T}}_{inter}^{10000}$ on the aforementioned datasets. ${\mathcal{T}}_{intra}^{0.5}$ displays lower runtimes than $Base$ across the datasets while ${\mathcal{T}}_{inter}^{100}$, ${\mathcal{T}}_{inter}^{1000}$, and ${\mathcal{T}}_{inter}^{10000}$ do so on COCO and NoCaps. We have shown how VWTs are more suitable for online inference by striking a good balance between savings in wattage and costs to runtime.
Dataset Model ${\mathcal{T}}_{inter}^{100}$ ${\mathcal{T}}_{inter}^{1000}$ ${\mathcal{T}}_{inter}^{10000}$
@@
OpenImages (Com.) - 171 - 171 - 176
OpenImages (Rare) - 169 - 169 - 174
COCO BLIP 405 406 409 408 442 440
@@
Table 4: Token length per sample (including [CLS]). ${\mathcal{T}}_{inter}^{{\mathcal{V}}}$ of varying pre-processing data and vocabulary sizes are shown. Visual words are formed from patches after the pre-trained convolution layer of CLIP or BLIP. Unlike Table 1, poor compression is seen due to dissimilarity between patches and new visual words.
4.4 VWTs and Visual Performance
Another important factor in compression is the effect on model performance. In Table 3, we tabulate the performance of image classification and captioning using CLIP and BLIP, respectively. For visual recognition and subgroup robustness, we analyze performance from an average and worst-group perspective as done by Sagawa et al. [50] and Romiti et al. [49]. First, we find the degradation in average accuracy to be small for the VWTs. The largest drop of up to 2% only is shown by ${\mathcal{T}}_{intra}^{0.5}$ on MetaShift. Likewise, both $Q_{8}$ and $ToME$ retain a high average accuracy across the datasets. Second, we observe possible improvements in subgroup robustness with VWTs. On Waterbirds and CelebA, the worst-group accuracy (WGA) with ${\mathcal{T}}_{intra}^{0.5}$ increases by up to 29% and 8%, respectively. Only on MetaShift do we observe a lower WGA than $Q_{8}$ or $ToME$ relative to $Base$. Like Gee et al. [20], we find compression to not always be harmful to subgroup robustness by improving the WGA at a negligible cost to overall performance.
To further validate performance, we conduct additional evaluations on the large-scale OpenImages v6 dataset for zero-shot classification. We report mean Average Precision (mAP) as done by Huang et al. [25]. Likewise, from Tables 2 and 3, we find the VWTs to be more suitable for online inference than $Q_{8}$ or $ToME$ in the large-scale setting while retaining comparable performance to $Base$.
For image captioning with BLIP, we evaluate our models following the setting in Li et al. [30] by using the BLEU, CIDEr, and SPICE metrics w.r.t. the ground truth captions. On COCO, the VWTs display a higher performance than $ToME$ and are competitive with $Q_{8}$. On NoCaps, we see the largest degradation on the out-of-domain samples with ${\mathcal{T}}_{inter}^{100}$. However, overall performance of ${\mathcal{T}}_{intra}^{0.5}$, ${\mathcal{T}}_{inter}^{1000}$, and ${\mathcal{T}}_{inter}^{10000}$ are still higher than $ToME$. Only $Q_{8}$ displays a slight improvement over $Base$. Like $ToME$, the VWTs are shown to not require additional end-to-end fine-tuning for performance retention.
The generated captions by BLIP [30] can be used as priors for further training of other models. As such, longer descriptive captions may be more desirable than the typical short captions associated with Internet data. In Figure 5, we visualize the long-form captions on COCO. To enable longer generations, we set the length penalty to 2.0 and double the maximum length to 40. All other generation settings are kept the same. With longer captions, the generation may degenerate into unnecessary repetitions on certain samples. Interestingly, descriptiveness and coherence improves more with ${\mathcal{T}}_{inter}^{100}$ than ${\mathcal{T}}_{intra}^{0.5}$, ${\mathcal{T}}_{inter}^{1000}$, or ${\mathcal{T}}_{inter}^{10000}$ inspite of its higher sequence compression as seen on COCO in Table 1.
Dataset Model ${\mathcal{T}}_{inter}^{100}$ ${\mathcal{T}}_{inter}^{1000}$ ${\mathcal{T}}_{inter}^{10000}$